# Beam Cross Sections Create Mixtures: Improving Feature Localization in Secondary Electron Imaging

### v1

@@ Abstract
Secondary electron (SE) imaging techniques, such as scanning electron microscopy and helium ion microscopy (HIM), use electrons emitted by a sample in response to a focused beam of charged particles incident at a grid of raster scan positions. Spot size—the diameter of the incident beam’s spatial profile—is one of the limiting factors for resolution, along with various sources of noise in the SE signal. The effect of the beam spatial profile is commonly understood as convolutional. We show that under a simple and plausible physical abstraction for the beam, though convolution describes the mean of the SE counts, the full distribution of SE counts is a mixture. We demonstrate that this more detailed modeling can enable resolution improvements over conventional estimators through a stylized application in semiconductor inspection of localizing the edge in a two-valued sample. We derive Fisher information about edge location in conventional and time-resolved measurements (TRM) and also derive the maximum likelihood estimate (MLE) from the latter. Empirically, the MLE computed from TRM is approximately efficient except at very low beam diameter, so Fisher information comparisons are predictive of performance and can be used to optimize the beam diameter relative to the raster scan spacing. Monte Carlo simulations show that the MLE gives a 5-fold reduction in root mean-squared error (RMSE) of edge localization as compared to conventional interpolation-based estimation. Applied to three real HIM datasets, the average RMSE reduction factor is 5.4.
Index Terms:
electron beams, electron microscopy, Fisher information, focused ion beams, mixture models, particle beam microscopy, Poisson compounding, semiconductor inspection, spot size effect, super-resolution localization, zero-truncated Poisson mixture.
@@ I Introduction
The micrographs produced by SE imaging instruments show spatial variation of SE yield $\eta$—defined as the mean number of SEs emitted by the sample per incident particle—subject to various sources of noise. Though the dependence of SE yield on local sample properties is complicated, SE imaging is nevertheless ubiquitous in semiconductor metrology, where it is used for critical dimension measurements [6], line width roughness analysis [7, 8, 9, 10, 11, 12], characterizing cross-sectional samples [13, 14], and analyzing and classifying sample defects [15, 16, 17]. Biological applications of SE imaging have been abundant for a half century [18] and are especially popular when combined with focused ion beam (FIB) sectioning for volumetric imaging [19, 20, 21].
The noise in SE images comes from three mechanisms [22, 23]: source shot noise, variation in the number of incident particles in a fixed time period; target shot noise, variation in the number of emitted SEs for a given incident particle; and detector noise, variation in measured response to a number of SEs (including the efficiency in emitted SEs reaching the detector). Additionally, the spot size or cross-sectional profile of the incident particle beam is not generally considered a source of noise but it too influences the image quality.
@@
In this paper, we complement the framework of previous works with detailed physics-based probabilistic modeling of the beam cross section. Our main assertion is that nonzero beam cross-sectional area causes the SE count due to a single incident particle to follow a mixture distribution. This contrasts with a convolutional model that is intuitive because, naïvely, the effect of the beam seems analogous to blurring in a linear optical system. Indeed, convolutional modeling is conventional in scanning electron microscopy [28, 29, 6]; we seek to expose the limitations and suboptimality of this modeling while providing an operational alternative.
@@
As a prototypical problem in semiconductor metrology, we consider here the localization of an axis-aligned straight edge between SE yield levels in a two-dimensional sample. This simplified imaging-based inspection problem allows us to use root mean-squared error (RMSE) to explicitly compare the accuracies of estimates produced from mixture modeling and from convolutional modeling. Monte Carlo simulations show RMSE reduced by a factor of 2.5 by using TRM with a mixture model as compared to TRM with a convolution model; improvement relative a conventional interpolation method is by a factor of 5. On experimental HIM data, we observe an an average RMSE reduction factor of 1.5 with respect to a convolution model and 5.4 with respect to interpolation.
@@
The paper is organized as follows. Section II introduces models of the beam, the sample, and their interaction that underlie this work. We establish there a physical justification for a two-component Poisson mixture model; the estimation advantage presented by this model motivates our entire study. Analyses in Section II are for observations of SE counts due to a single incident particle and for deterministic repetitions of such hypothetical observation trials. In practice, particle incidences are random and not perfectly inferrable. Mitigating this effect was the motivation for time-resolved measurement. Section III reviews—and in some cases slightly generalizes—results from [25, 26] on the advantages from TRM when beam cross-sectional effects are neglected. Section IV brings together the spatial (Section II) and temporal (Section III) aspects of our incident beam model. Section V analyzes estimation of edge location based on the mixture model when combining data acquired from a line of scan locations across an edge. We show that accuracy of localization can be finer than scan grid spacing, thereby achieving super-resolution. We furthermore establish a method for optimal selection of beam width. In Monte Carlo simulations of estimator performance, we observe that the mixture estimator has lower mean-squared error (MSE) than other estimators of the edge location. The same performance trends are also observed on experimental datasets acquired on a helium ion microscope.
@@ II-A Model of Incident Particle Beam
As shown in the cartoon of Fig. 1(a), incident particles are haphazard in both transverse spatial positions and times of incidence. We model the beam temporally as arrivals in a Poisson process with a known rate $\Lambda$ per unit time (see Fig. 1(b)). Let $t$ denote the pixel dwell time. Then the number of particles $M$ incident on a pixel is a Poisson random variable with parameter $\lambda=\Lambda t$. We refer to $\lambda$ as the dose; normalization by beam cross-sectional area or raster scan density is helpful when assessing likelihood of sample damage. Recall that $M$ is not directly observable; its variation from $\mathrm{E}[\,{M}\,]=\lambda$ is called source shot noise.
@@ II-B Model of Sample
We normalize the transverse spatial dimensions of the sample by the imaging raster scan spacing so that raster scan locations are on the grid $\{0,\,1,\,\ldots,\,\ell-1\}^{2}$. Our mathematical model of the sample is not restricted to the raster scan grid because we wish to model effects of the beam cross section and the potential for sub-pixel resolution. Thus, the sample is defined as a function $\eta:{\mathbb{R}}^{2}\rightarrow[0,\infty)$ that maps from the ion incidence position (which, due to nonzero beam width, is not exactly a raster scan position) to the SE yield.
When an ion is incident at $\mathbf{s}=(s_{1},s_{2})$, the number of emitted SEs is not deterministically $\eta(\mathbf{s})$. Distinct mechanisms of emission have been established. Potential emission depends mostly on incident particle charge and results in SE counts that can be modeled as binomial [35, 36]; it is especially significant with multiply-charged projectiles [37, 38]. Kinetic emission depends mostly on incident particle mass and velocity and results in SE counts that can be modeled as Poisson when cascade effects are minimal and Pólya more generally [39]. Due to the combination of mechanisms and the crudeness of the theoretical models, no single parametric model fits measured data well across a wide range of conditions. Many studies of SE emission and all previous SE imaging methods, to the best of our knowledge, treat the number of SEs as a Poisson random variable, including effects of nonzero beam width. Here, we assume the SE counts are Poisson distributed when effects of nonzero beam width are not included.
The problem of SE imaging in general is the estimation of the function $\eta$, usually on a subset of the domain such as $0,\,\ell-1]^{2}$ or only on the scan grid $\{0,\,1,\,\ldots,\,\ell-1\}^{2}$. In Section II-D, we will restrict $\eta$ to model a two-valued sample with a single axis-aligned edge.
II-C Models of SE Distribution for One Incident Particle
Under the beam and sample models in Sections II-A and II-B, one incident ion at grid scan location $\mathbf{g}=(g_{1},g_{2})$ will actually strike the sample at $\mathbf{s}=\mathbf{g}+\mathbf{W}$, where $\mathbf{W}\sim{\mathcal{N}(\mathbf{0},\sigma_{\mathrm{b}}^{2}I_{2})}$. Since the sample has SE yield $\eta(\mathbf{s})$ for this incidence position, the random number of generated SEs is
$X\sim\mathop{\mathrm{Poisson}}(\eta(\mathbf{g}+\mathbf{W})).$ (1)
Since the random quantity $\eta(\mathbf{g}+\mathbf{W})$ is the parameter of a Poisson random variable, this $X$ has a Poisson mixture distribution [40].
@@ II-C Models of SE Distribution for One Incident Particle
$\mathrm{E}\!\left[\,{X}\,\right]$ $=\mathrm{E}\!\left[\,{\mathrm{E}\!\left[\,{X\,|\,\mathbf{W}}\,\right]}\,\right]=\mathrm{E}\!\left[\,{\eta(\mathbf{g}+\mathbf{W})}\,\right]$
$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\eta(\mathbf{g}+\mathbf{w})f_{\mathbf{W}}(w_{1},w_{2})\,\mathsf{d}w_{1}\,\mathsf{d}w_{2}.$ (2)
The convolutional form in (2) is reminiscent of a linear blur kernel in an optical system. It is important, however, to make the distinction between describing the full distribution or merely the mean. Since SE distributions are often assumed to be Poisson, we will refer to
$X_{\rm convolution}\sim\mathop{\mathrm{Poisson}}(\mathrm{E}\!\left[\,{\eta(\mathbf{g}+\mathbf{W})}\,\right])$ (3)
as the convolutional model for $X$.
The mixture and convolutional models are defined to match in mean. A simple way to illustrate that they are not equal in distribution is to compute variances. Since the convolutional model is Poisson distributed, its variance is plainly
$\mathrm{var}\!\left({X_{\rm convolution}}\right)=\mathrm{E}\!\left[\,{\eta(\mathbf{g}+\mathbf{W})}\,\right].$ (4)
For the mixture model, the variance is straightforward using the law of total variance:
$\mathrm{var}\!\left({X}\right)$ $=\mathrm{E}\!\left[\,{\mathrm{var}\!\left({X\,|\,W}\right)}\,\right]+\mathrm{var}\!\left({\mathrm{E}\!\left[\,{X\,|\,W}\,\right]}\right)$
$=\mathrm{E}\!\left[\,{\eta(\mathbf{g}+\mathbf{W})}\,\right]+\mathrm{var}\!\left({\eta(\mathbf{g}+\mathbf{W})}\right).$ (5)
We will refer to
$\mathrm{var}\!\left({X}\right)-\mathrm{var}\!\left({X_{\rm convolution}}\right)=\mathrm{var}\!\left({\eta(\mathbf{g}+\mathbf{W})}\right)$ (6)
as the excess variance of the mixture model. It is nonnegative, and it is zero only in the limit of $\eta$ have no variation in the neighborhood of the nominal scan location $\mathbf{g}$, where the relevant neighborhood size depends on $\sigma_{\mathrm{b}}$.
We will reinterpret (3)–(6) after restricting our attention to samples with a single straight edge (Section II-D). With that restriction, we will examine the estimability of $\eta$ under the two models (Section II-E).
@@ II-D Samples with Two SE Yield Values
as illustrated in Fig. 2(a). A step in semiconductor inspection or manufacturing process characterization could include imaging of this sample. In the case that $\eta_{1}$ and $\eta_{2}$ are known, this imaging task reduces to estimation of edge position $\gamma$.
@@
Figure 2: A sample with two SE yield values separated by a vertical edge at horizontal position $\gamma$. (a) Depiction of the function $\eta(s_{1},s_{2})$. (b) Sample and beam distribution projected to 1D functions of only the the horizontal position.
Since the sample has no vertical variation and the Gaussian beam model in Section II-A has independent horizontal and vertical deviations from the raster scan position, the number of generated SEs $X$ depends only on the horizontal raster scan position. Specifically, from (1), the distribution of $\eta(\mathbf{g}+\mathbf{W})$ determines the distribution of $X$. Since (7) has no dependence on $s_{2}$ and $W_{1}$ is independent of $W_{2}$, we are left with projections to 1D functions as shown in Fig. 2(b). The projected beam profile is Gaussian with standard deviation $\sigma_{\mathrm{b}}$.
The two-valued nature of $\eta$ allows us to simplify further. Define
$q=$ $\ \mathrm{P}({\eta(\mathbf{g}+\mathbf{W})=\eta_{2}})$
@@
where (a) follows from (7); and (b) uses $\Phi(\cdot)$ to denote the standard normal cumulative distribution function. This $q$ is the mixing parameter in a two-component Poisson mixture distribution for $X$. Since $\Phi$ is one-to-one, an estimate of $q$ can be mapped to an estimate of edge position $\gamma$ using (8).
For our two-valued sample, we can revisit the difference between the mixture and convolutional models introduced in Section II-C. The mixture model is
@@
which has mean
@@
The convolutional model—which we posit to be inaccurate despite matching in mean—is
$X_{\rm convolution}\sim\mathop{\mathrm{Poisson}}((1-q)\eta_{1}+q\eta_{2}).$ (12)
@@
$\mathrm{var}\!\left({X}\right)-\mathrm{var}\!\left({X_{\rm convolution}}\right)=q(1-q)(\eta_{2}-\eta_{1})^{2}.$ (13)
II-E SE Yield Estimation with Deterministic Number of Ions
Let us now consider how the mixture model improves the ability to estimate SE yield. Suppose one is able to observe SE counts $X_{1},\,X_{2},\,\ldots,\,X_{m}$ from $m$ incident ions. In the absence of a sophisticated model, one would conventionally estimate the mean $\mathrm{E}\!\left[\,{X}\,\right]$ using the sample mean of the observations,
$\widehat{\eta}_{\mathrm{conv}}=\frac{1}{m}\sum_{i=1}^{m}X_{i}.$ (14)

### v2

@@ Abstract
Secondary electron (SE) imaging techniques, such as scanning electron microscopy and helium ion microscopy (HIM), use electrons emitted by a sample in response to a focused beam of charged particles incident at a grid of raster scan positions. Spot size—the diameter of the incident beam’s spatial profile—is one of the limiting factors for resolution, along with various sources of noise in the SE signal. The effect of the beam spatial profile is commonly understood as convolutional. We show that under a simple and plausible physical abstraction for the beam, though convolution describes the mean of the SE counts, the full distribution of SE counts is a mixture. We demonstrate that this more detailed modeling can enable resolution improvements over conventional estimators through a stylized application inspired by semiconductor inspection: localizing the edge in a two-valued sample. We derive Fisher information about edge location in conventional and time-resolved measurements (TRM) and also derive the maximum likelihood estimate (MLE) from the latter. Empirically, the MLE computed from TRM is approximately efficient except at very low beam diameter, so Fisher information comparisons are predictive of performance and can be used to optimize the beam diameter relative to the raster scan spacing. Monte Carlo simulations provide an example of the MLE giving a 5-fold reduction in root mean-squared error (RMSE) of edge localization as compared to conventional interpolation-based estimation. The RMSE is substantially below both the beam diameter and the raster scan spacing and thus sub-pixel localization is demonstrated. Applied to three real HIM datasets, the average RMSE reduction factor is 5.4.
@@ I Introduction
The micrographs produced by SE imaging instruments show spatial variation of SE yield—defined as the mean number of SEs emitted by the sample per incident particle—subject to various sources of noise. Though the dependence of SE yield on local sample properties is complicated, SE imaging is nevertheless ubiquitous in semiconductor metrology, where it is used for critical dimension measurements [6], line width roughness analysis [7, 8, 9, 10, 11, 12], characterizing cross-sectional samples [13, 14], and analyzing and classifying sample defects [15, 16, 17]. Biological applications of SE imaging have been abundant for a half century [18] and are especially popular when combined with focused ion beam (FIB) sectioning for volumetric imaging [19, 20, 21].
The noise in SE images comes from three mechanisms [22, 23]: source shot noise, variation in the number of incident particles in a fixed time period; target shot noise, variation in the number of emitted SEs for a single incident particle; and detector noise, variation in measured response to a number of SEs (including the efficiency in emitted SEs reaching the detector). Additionally, the spot size or cross-sectional profile of the incident particle beam is not generally considered a source of noise but it too influences the image quality.
@@
In this paper, we complement the framework of previous works with detailed physics-based probabilistic modeling of the beam cross section. Our main assertion is that nonzero beam cross-sectional area causes the SE count due to a single incident particle to follow a mixture distribution. This contrasts with a convolutional model that may be intuitive because, naïvely, the effect of the beam seems analogous to blurring in a linear optical system. Indeed, convolutional modeling is conventional in scanning electron microscopy [28, 29, 6]; we seek to expose the limitations and suboptimality of this modeling while providing an operational alternative.
@@
As a simplified abstraction of a prototypical problem in semiconductor metrology, we consider here the localization of an axis-aligned straight edge in a two-dimensional sample with distinct SE yield levels on each side of the edge. This simplified imaging-based inspection problem allows us to use root mean-squared error (RMSE) of the edge location to explicitly compare the accuracies of estimates produced from mixture modeling and from convolutional modeling. In a highlighted example (see Fig. 8(b)), Monte Carlo simulations show RMSE reduced by a factor of 2.5 by using TRM with a mixture model as compared to TRM with a convolution model; improvement relative to a conventional interpolation method is by a factor of 5. In general, improvements will depend on many factors, including beam diameter relative to raster scan spacing, edge location relative to the scan grid, SE yield values, and number of incident particles. On experimental HIM data, we observe an average RMSE reduction factor of 1.5 with respect to a convolution model and 5.4 with respect to interpolation.
@@
Advantages from our novel mixture modeling are small in the absence of TRM, which may explain why this type of modeling has not been developed previously. Here, this dependency necessitates detailed discussion of spatial and temporal beam characteristics and the properties of TRM itself before we turn to the application of edge localization. Section II introduces the models of the beam (Section II-A), the sample (Section II-B), and their interaction (Section II-C) that underlie this work. We establish there a physical justification for a two-component Poisson mixture model for SE counts when the sample is two-valued with SE yields $\eta_{1}$ and $\eta_{2}$ (Section II-D). Though we are not yet considering the edge localization problem per se—and for clarity we do not include source shot noise—the fundamental gap between the best uses of TRM and non-TRM data when $\eta_{1}$ and $\eta_{2}$ are known (Section II-E) motivates our entire study. The challenge is to realize the potential gains from a mixture model, translated to edge localization, in the face of source shot noise (Section II-F). Analyses in Section II are for observations of SE counts due to a single incident particle and for repetitions of such hypothetical observations for a deterministic number of trials. In practice, particle incidences are random and not perfectly inferrable. Mitigating this effect was the motivation for time-resolved measurement. Section III reviews—and in some cases slightly generalizes—results from [25, 26] on the advantages from TRM when beam cross-sectional effects are neglected. Measurement models are formally introduced (Section III-A), and analyses are provided for conventional (Section III-B) and time-resolved measurements (Section III-C). Section IV brings together the spatial (Section II) and temporal (Section III) aspects of our incident beam model for the estimation of two-component Poisson mixtures, for conventional (Section IV-A) and time-resolved measurements (Section IV-B). Using the relationship between edge location, mixture fraction, and SE yield, Section V analyzes estimation of edge location when combining data acquired from a line of scan locations across an edge. We show that accuracy of localization can be finer than scan grid spacing, thereby achieving sub-pixel resolution. We furthermore establish a method for optimal selection of beam width. In Monte Carlo simulations of estimator performance, we observe that the mixture estimator has lower mean-squared error (MSE) than other estimators of the edge location. The same performance trends are also observed on experimental datasets acquired on a helium ion microscope.
@@ II-A Model of Incident Particle Beam
As shown in the cartoon of Fig. 1(a), incident particles are haphazard in both transverse spatial positions and times of incidence. We model the beam temporally as arrivals in a Poisson process with a known rate $\Lambda$ per unit time (see Fig. 1(b)). Let $t$ denote the pixel dwell time. Then the number of particles $M$ incident on a pixel is a Poisson random variable with parameter $\lambda=\Lambda t$. We refer to $\lambda$ as the dose; normalization by beam cross-sectional area or raster scan density is helpful when assessing likelihood of sample damage. The value of $M$ is not directly observable; its variation from $\mathrm{E}[\,{M}\,]=\lambda$ is called source shot noise.
@@
In an ion- or electron-beam instrument, the values of beam intensity $\Lambda$ and beam standard deviation $\sigma_{\mathrm{b}}$ depend on many factors and are controllable to some extent, with limitations that depend on the particle type. Increasing $\Lambda$ is generally associated with increasing $\sigma_{\mathrm{b}}$. Increasing working distance is also associated with higher $\sigma_{\mathrm{b}}$. Typical minimum values of $\sigma_{\mathrm{b}}$ are in the subnanometer range, but larger values are chosen (e.g., through beam defocus) to limit sample damage. One contribution of this paper is to study favorable ratios between $\sigma_{\mathrm{b}}$ and raster scan spacing.
@@ II-B Model of Sample
We normalize the transverse spatial dimensions of the sample by the imaging raster scan spacing so that raster scan locations are on the grid $\{0,\,1,\,\ldots,\,\ell-1\}^{2}$. Our mathematical model of the sample is not restricted to the raster scan grid because we wish to model effects of the beam cross section and the potential for sub-pixel resolution. Thus, the sample is defined as a function $\varphi:{\mathbb{R}}^{2}\rightarrow[0,\infty)$ that maps from the ion incidence position (which, due to nonzero beam width, is not exactly a raster scan position) to the SE yield.
When an ion is incident at $\mathbf{s}=(s_{1},s_{2})$, the number of emitted SEs is not deterministically $\varphi(\mathbf{s})$. Distinct mechanisms of emission have been established. Potential emission depends mostly on incident particle charge and results in SE counts that can be modeled as binomial [35, 36]; it is especially significant with multiply-charged projectiles [37, 38]. Kinetic emission depends mostly on incident particle mass and velocity and results in SE counts that can be modeled as Poisson when cascade effects are minimal and Pólya more generally [39]. Due to the combination of mechanisms and the crudeness of the theoretical models, no single parametric model fits measured data well across a wide range of conditions. Many studies of SE emission and all previous SE imaging methods, to the best of our knowledge, treat the number of SEs as a Poisson random variable, including effects of nonzero beam width. Here, we assume the SE counts are Poisson distributed when effects of nonzero beam width are not included.
The problem of SE imaging in general is the estimation of the function $\varphi$, usually on a subset of the domain such as $0,\,\ell-1]^{2}$ or only on the scan grid $\{0,\,1,\,\ldots,\,\ell-1\}^{2}$. In Section II-D, we will restrict $\varphi$ to model a two-valued sample with a single axis-aligned edge.
II-C Model of SE Distribution for One Incident Particle
Under the beam and sample models in Sections II-A and II-B, one incident ion at grid scan location $\mathbf{g}=(g_{1},g_{2})$ will actually strike the sample at $\mathbf{s}=\mathbf{g}+\mathbf{W}$, where $\mathbf{W}\sim{\mathcal{N}(\mathbf{0},\sigma_{\mathrm{b}}^{2}I_{2})}$. Since the sample has SE yield $\varphi(\mathbf{s})$ for this incidence position, the random number of generated SEs is
$X\sim\mathop{\mathrm{Poisson}}(\varphi(\mathbf{g}+\mathbf{W})).$ (1)
Since the random quantity $\varphi(\mathbf{g}+\mathbf{W})$ is the parameter of a Poisson random variable, this $X$ has a Poisson mixture distribution [40].
@@ II-C Model of SE Distribution for One Incident Particle
$\mathrm{E}\!\left[\,{X}\,\right]$ $=\mathrm{E}\!\left[\,{\mathrm{E}\!\left[\,{X\,|\,\mathbf{W}}\,\right]}\,\right]=\mathrm{E}\!\left[\,{\varphi(\mathbf{g}+\mathbf{W})}\,\right]$
$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\varphi(\mathbf{g}+\mathbf{w})f_{\mathbf{W}}(w_{1},w_{2})\,\mathsf{d}w_{1}\,\mathsf{d}w_{2}.$ (2)
We use $\eta$ as a compact shorthand for $\mathrm{E}\!\left[\,{X}\,\right]$, which is the SE yield incorporating the beam model, in contrast to the idealized $\varphi(\mathbf{g})$. The convolutional form in (2) is reminiscent of a linear blur kernel in an optical system. It is, however, central to our premise to describe the full distribution rather than merely the mean.
A simple way to illustrate that $X$ is not Poisson distributed—and hence the effect of the beam cross-section is not merely a convolutional blurring of $\varphi(\mathbf{s})$, decoupled from the distribution family of the target shot noise—is to compute its variance. Using the law of total variance:
$\mathrm{var}\!\left({X}\right)$ $=\mathrm{E}\!\left[\,{\mathrm{var}\!\left({X\,|\,\mathbf{W}}\right)}\,\right]+\mathrm{var}\!\left({\mathrm{E}\!\left[\,{X\,|\,\mathbf{W}}\,\right]}\right)$
$=\mathrm{E}\!\left[\,{\varphi(\mathbf{g}+\mathbf{W})}\,\right]+\mathrm{var}\!\left({\varphi(\mathbf{g}+\mathbf{W})}\right).$ (3)
Since Poisson-distributed $X$ would have variance equal to its mean, we will refer to
$\mathrm{var}\!\left({X}\right)-\mathrm{E}\!\left[\,{X}\,\right]=\mathrm{var}\!\left({\varphi(\mathbf{g}+\mathbf{W})}\right)$ (4)
as the excess variance of the mixture model. It is nonnegative, and it is zero only in the limit of $\varphi$ having no variation in the neighborhood of the nominal scan location $\mathbf{g}$, where the relevant neighborhood size depends on $\sigma_{\mathrm{b}}$.
@@ II-D Samples with Two SE Yield Values
as illustrated in Fig. 2(a). A step in semiconductor inspection or manufacturing process characterization could include imaging of this sample. Except when discussing the value of a priori knowledge of $\eta_{1}$ and $\eta_{2}$ in Section II-E, we will assume $\eta_{1}$ and $\eta_{2}$ are known, with $\eta_{1}<\eta_{2}$. The imaging task then reduces to estimation of edge position $\gamma$.
@@
Figure 2: A sample with two SE yield values separated by a vertical edge at horizontal position $\gamma$. (a) Depiction of the function $\varphi(s_{1},s_{2})$. (b) Sample and beam distribution projected to 1D functions of only the horizontal position.
Since the sample has no vertical variation and the Gaussian beam model in Section II-A has independent horizontal and vertical deviations from the raster scan position, the number of generated SEs $X$ depends only on the horizontal raster scan position. Specifically, from (1), the distribution of $\varphi(\mathbf{g}+\mathbf{W})$ determines the distribution of $X$. Since (5) has no dependence on $s_{2}$ and $W_{1}$ is independent of $W_{2}$, we are left with projections to 1D functions as shown in Fig. 2(b). The projected beam profile is Gaussian with standard deviation $\sigma_{\mathrm{b}}$.
The two-valued nature of $\varphi$ allows us to simplify further. Define
$q=$ $\ \mathrm{P}({\varphi(\mathbf{g}+\mathbf{W})=\eta_{2}})$
@@
where (a) follows from (5); and (b) uses $\Phi(\cdot)$ to denote the standard normal cumulative distribution function. Using this $q$ in (1) makes $X$ a two-component Poisson mixture:
@@
meaning that the probability mass function (PMF) of $X$ is
Since $\Phi$ is one-to-one, an estimate of $q$ can be mapped to an estimate of edge position $\gamma$ using (6).
Using the law of iterated expectation, the mean of $X$ is
@@
$\mathrm{var}\!\left({X}\right)-\mathrm{E}\!\left[\,{X}\,\right]=q(1-q)(\eta_{2}-\eta_{1})^{2}.$ (11)
II-E Potential Improvement from Mixture Model
Let us now demonstrate that exploiting the mixture model has the potential to improve the ability to estimate SE yield; this is translated to edge localization in Section V after we incorporate source shot noise. Suppose $m$ ions are incident, with ion $i$ resulting in $X_{i}$ observed SE counts, for $i=1,\,2,\,\ldots,\,m$. We will analyze estimators that do and do not use prior knowledge of $\eta_{1}$ and $\eta_{2}$ in the estimation of $\eta=\mathrm{E}\!\left[\,{X}\,\right]$.
When $q$, $\eta_{1}$, and $\eta_{2}$ are all unknown during estimation, one would naturally estimate $\eta$ using the sample mean of the observations,
$\widehat{\eta}_{\mathrm{SM}}=\frac{1}{m}\sum_{i=1}^{m}X_{i}.$ (12)